Find two consecutive whole numbers such that the sum of their squares is 313.
Answer by flame8855(424)

x^2 + y^2 = 313
y=x+1
x^2 +x^2 +1 +2x = 313
2x^2 +2x-312 = 0
x^2 +x -156= 0
x = 12
y = 13